(3 pts) Suppose y, the number of cases of a disease, is reduced by 12% per year.
(a) If there are initially 10,000 cases, express y as a function of t, the number of years elapsed.
y = (do not enter any commas in your formula)
(b) How many cases will there be 7 years from now?
cases.
(c) How long does it take to reduce the number of cases to 1000?
years

Answer by ewatrrr(24785)

Q(t) = Q(0)e^(-kt)
Q(t) = 10000e^(-.12t)
Q(7) = 10000e^(-.12*7)
|(c) How long does it take to reduce the number of cases to 1000?
1000/10000 = e^(-.12*t)
ln(.1) = -.12t
ln(.10)/.12 = t